I am stuck on this question and don't know where to start.
The width of a rectangle is 10 mm shorter than its length. Its perimeter is 244 mm. Let l equal the length of the rectangle.
1. Write an expression for the width in terms of the length.
2. Write an equation for the perimeter in terms of l.
3. Solve the equation, clearly indicating the length of the rectangle.
4. Solve for the width of the rectangle. Clearly indicate the width of the rectangle.

Answer by josgarithmetic(39617)

Let WHAT equal the length of the rectangle?
Use this little change. Let L equal the length of the rectangle.

If width is 10 mm shorter than L, then this means L is 10 mm more than the width. Either way, you have these:

Length is L
Width is L-10
Also given is perimeter. 2(L+w)=244, using w for width.
Putting in all the expressions, 2(L+(L-10))=244, and you want to find the value for L.

2(L+(L-10))=244
2(2L-10)=244
2L-10=122
L-5=61
L=66---------part of the result, the main important result for the solution